OpenStudy (anonymous):

find the sum of series 1*2+2*3+3*4+. . .152*153+153*154

OpenStudy (anonymous):

\[\Huge \sum_{n=1}^{153}n(n+1)\]

OpenStudy (anonymous):

Thanks got to that part. How do u sum that expression?

OpenStudy (anonymous):

Use distributive property.

OpenStudy (anonymous):

\[\Huge \sum_{n=1}^{153}n^2 + n = \sum_{n=1}^{153} n^2 + \sum_{n=1}^{153}n\]

OpenStudy (anonymous):

Thanks that gives \[(\frac{ 1 }{ 6} N (2N+1)+(N+1))+(\frac{ 1 }{ 2}N(n+1))\]